# Index Terms: smartphone, vibration, tactile, perceptionI Introduction

### 14 Jan 2024 v1

@@ Abstract
Tactile perception plays an important role in activities of daily living, and it can be impaired in individuals with medical conditions. The most common tools used to assess tactile sensation, the Semmes-Weinstein monofilaments and the 128 Hz tuning fork, have poor repeatability and resolution. Long term, we aim to provide a repeatable, high-resolution testing platform that can be used to assess vibrotactile perception through smartphones without the need for an experimenter to be present to conduct the test. We present a smartphone-based vibration perception measurement platform and compare its performance to measurements from standard monofilament and tuning fork tests. We conducted a user study with 36 healthy adults in which we tested each tool on the hand, wrist, and foot, to assess how well our smartphone-based vibration perception thresholds (VPTs) detect known trends obtained from standard tests. The smartphone platform detected statistically significant changes in VPT between the index finger and the foot and also between the feet of younger adults and older adults. Our smartphone-based VPT had a moderate correlation to tuning fork-based VPT. A long-term objective of this work is to develop an accessible smartphone-based platform that can be used to measure disease progression and regression.
@@
Tactile perception, including vibrotactile perception, plays a critical role in enabling humans to perform various sensorimotor tasks such as object manipulation, navigation, and playing sports [1, 2, 3, 4]. We even use vibrotactile perception to balance during walking, a pervasive activity of daily living [5, 6, 7, 8, 9]. Many underlying health conditions including diabetes, chemotherapy, and direct injuries to the body can impair our tactile perception [10]. Given our reliance on tactile perception, deficits in perceiving tactile cues can have devastating consequences. Assessing tactile perception can help us understand disease progression and recovery, especially in response to treatment.
@@
Frequent use of vibration perception testing occurs during routine screenings of people experiencing peripheral neuropathy. The most common clinical diagnostics for peripheral neuropathy are the Semmes-Weinstein monofilament exam and the 128 Hz tuning fork exam [11]. Although these exams provide important information on tactile ability, they have some limitations in the resolution of data they are able to provide. Tuning forks often provide inconsistent vibrations due to differences in how the clinician strikes the fork [12]. Monofilaments also suffer from variations in force delivered due to variations in clinician application and overuse [13]. Also, when used in many clinical settings, both of these tools only measure a binary response (‘yes, can feel’ or ‘no, cannot feel’) to a single provided stimulus.
@@
Given the rise of smartphones with high quality vibration actuators, there has been a rising interest in conducting mobile haptics experiments [14]. We conducted preliminary work which showed that smartphone vibrations are more repeatable than tuning fork vibrations and that vibration perception threshold could be reliably measured using a custom application employing a staircase algorithm [15]. Also in 2022, Torres et al. [16] characterized smartphone vibrations and found that smartphone-based vibration perception thresholds correlate with monofilament-based pressure thresholds in the index finger. Several researchers have also worked toward validating the use of smartphone vibrations for diagnosing diabetic peripheral neuropathy. May et al. [17] showed that vibrations generated from a mobile phone could be used to detect diabetic peripheral neuropathy, and that the most accurate testing location was the first metatarsal head (a bony prominence in the big toe). Jasmin et al. [18] found that a vibration-based smartphone application has a moderate to strong correlation to tuning forks in classifying participants as experiencing neuropathy or not experiencing neuropathy, and that the interrater reliability using the smartphone application is high. While these studies are important steps towards developing an improved measurement tool, they suffer from various limitations including confounding factors in the study design, non-autonomous smartphone vibration perception data collection (reliance on an experimenter to physically conduct the exam), and the use of a binary measurement (yes/no) as opposed to a continuous, numerical value.
@@
In this work, we build upon our preliminary work by testing the feasibility of an application designed to measure smartphone-based vibration perception thresholds at different sites on the body (hand, wrist, and foot) in a healthy, age-diverse population. We also assess how the smartphone-based vibration perception thresholds (VPTs) correlate to both tuning fork vibration perception thresholds and monofilament force perception thresholds at these locations. Our main goal is to confirm that smartphones can be used for vibration perception measurements and capture similar trends as currently used methods. In Section II, we describe our methods in characterizing, designing, and administering the smartphone-based VPT, performing the tuning fork and monofilament exams, and conducting a user study and accompanying analyses. We then discuss the study results in Section III and key takeaways and future work, such as testing our smartphone application on various patient populations and expanding this application to additional types of smartphones, in Section IV.
@@ II-A1 Smartphone VPT Exam
We developed an iOS application that controls Apple’s Core Haptics parameters (“hapticIntensity” and “hapticSharpness”) and autonomously implements a staircase algorithm (reversals = 8) to measure vibration perception threshold. We chose to implement a one-up/one-down staircase method because it is a simple and fast way to calculate an absolute threshold. The one-up/one-down staircase method targets a performance level of 50% [19, 20], which is the standard performance level for absolute thresholds [19]. We chose to implement a one-up/one-down staircase instead of alternative one-up/two-down or one-up/three-down staircase methods that target higher performance levels because these alternatives require more trials and therefore more time to calculate a threshold; we chose to prioritize time as our eventual goal is to use this tool in a clinical setting where time is limited. Prior work by Yoshida and Kiernan, et al. [21] characterized the acceleration outputs at various locations on the phone; their results indicate that the output accelerations occur at a constant frequency of 230 Hz and that there is a positive, nonlinear relationship between the amplitude of the output acceleration and commanded “hapticIntensity” value when “hapticSharpness” is held constant at 1 and “hapticIntensity” is varied between 0.1 and 0.3. We further characterized this nonlinear relationship when “hapticSharpness” is held at a constant 1 and “hapticIntensity” is varied between 0 and 1 (Fig. 1).
@@
Despite this minor nonlinearity, we designed our staircase algorithm to output a continuous vibration for 0.1 seconds with the “hapticSharpness” set at 1.0 and the “hapticIntensity” varying with each step. This results in repeatable vibrations of varying amplitude with a constant frequency of 230 Hz which the users can respond to quickly and easily (Fig. 1). We chose to use “hapticSharpness” = 1 (230 Hz) since the Apple iPhone XS Max (our phone model) uses a linear resonant actuator that is tuned to operate at “hapticSharpness” = 1 (230 Hz). In addition to 230 Hz being the resonant frequency for this phone model’s linear resonant actuator, this frequency is also well within the range of the Pacinian corpuscles that are largely responsible for our ability to feel vibrations [22].
@@
The vibration amplitude starts small (“hapticIntensity” = 0.05) and increases with a step size of 0.05 until the vibration is detected by the user. At this point, the user says ‘yes’ to indicate detection, the spoken ‘yes’ is interpreted by the app using “Speech” (Apple’s voice recognition framework), and a reversal is recorded in the app. The vibration’s amplitude then decreases with a step size of 0.05 until the user can no longer detect the vibration (does not provide a spoken ‘yes’ response). Then, another reversal is recorded, the vibration’s amplitude increases again, and the staircasing of the “hapticIntensity” values continues until eight reversals are complete. Once complete, a CSV file containing the trial data is exported and stored in Google Firebase. The vibration perception threshold is calculated by averaging the “hapticIntensity” values at the eight reversal points (where the reversal point is the average of the value of the response that triggered a reversal and the value of the response prior to the reversal) as shown in Fig. 1. Time intervals between vibrations for each trial were randomly selected to reduce bias (ranging from 3-6 s), and responses had to occur within 2.5 seconds of the vibration in order to be recorded as a true positive response (as this is the upper end of haptic response times reported in literature [23, 24]). The smartphone VPT ranges from “hapticIntensity” of 0.05 (smallest vibration detected) to 1 (largest possible vibration output via Apple’s framework). As such, lower scores indicate better perception. If the largest vibrations output by the phone (vibrations generated by setting “hapticIntensity” to 1) were not felt three times in a row, the threshold for that trial is recorded as NaN because their threshold was outside the range of the smartphone-generated vibration stimuli. At each body part, smartphone vibration threshold is measured five times. The thresholds are then averaged, so that each participant’s average threshold at each of the body parts is reported.
@@
We measure vibration perception threshold at six locations of interest: the index finger pad, the back of the index finger, the pinky finger pad, the outer wrist, the inner wrist, and the big toe pad (Fig. 2). These locations were chosen both for both clinical relevance and ease of smartphone placement [25]. In Fig. 2B we also include the filtered vibration waveforms for the six configurations shown in Fig. 2A. For these measurements “hapticIntensity” was set to 0.25 and “hapticSharpness” was set to 1. We measure these waveforms using the setup described in detail in [15] and [21], with the modification that the phone was placed on a pillow while the body part was placed on the phone in the six different configurations. For each participant, smartphone-based VPT measurements are collected five times at each location, so that we can report the participant’s average smartphone-based VPT at each location.
@@
All smartphone-based measurements are collected on an Apple iPhone XS Max. Participants sit in a chair during the entire exam. For the finger and wrist locations, the phone is placed on a pillow that is placed on a desk. For the foot location, the phone is placed on a pillow, that is placed on the floor. To prevent participants from relying on hearing instead of feeling, to identify vibrations, participants wear headphones playing their preferred music. All but two participants listened to a Disney Hits playlist. Because the smartphone data collection component of the experiment took around an hour for most participants, we did not use white noise, which was found to make participants drowsy during piloting.
@@ II-A2 Tuning Fork Exam
We use a 128 Hz clinical tuning fork (CynaMed) to test the same six body parts as the smartphone (Fig. 2). Participants sit in a chair during the entire exam. For the finger and wrist locations, the hand/wrist is placed on a pillow that is placed on a desk. For the foot location, the foot is placed on a pillow, that is placed on a coffee table. The experimenter strikes the tines of the tuning fork on her knee and then places the base of the tuning fork on the body of the participant. Prior to collection, participants are touched with a vibrating tuning fork so they could get a sense of what it feels like and so that they understand that the sensation is not painful. Participants wear a blindfold so that they can not see when the tuning fork makes contact with their body, but do not wear headphones as the sound of the tuning fork vibrations are not easily discernible. Participants are instructed to say “start” when they start feeling a vibration and “stop” when they no longer feel any vibrations. A digital watch displaying seconds is used to measure the amount of time that the participant feels the vibration. The experimenter notes the second when the participate says “start” and also the second when the participant says “stop” and then records the difference in seconds, mimicking methods used in clinical settings. As such, higher tuning fork perception times indicate better perception. At each body part, tuning fork vibration perception time is measured five times. The times are then averaged, so that each participant’s average perception time at each of the body parts is reported.
@@ II-A3 Monofilament Exam
We use a 20-piece Semmes-Weinstein monofilament set (Touch Test Sensory Evaluators, North Coast Medical, Inc.) to assess light force perception. Filaments range from 0.008 grams-force to 300 grams-force. To assess light force perception at each of the six body parts (Fig. 2), we begin with the monofilament deemed normal for that location (0.07 grams-force for hands and dorsal feet, 3.61 grams-force for the plantar feet) [26]. Participants sit in a chair during the entire exam. For the finger and wrist locations, the hand/wrist is placed on a pillow that is placed on a desk. For the foot location, the foot is placed on a pillow, that is placed on a coffee table. Prior to collection, participants were touched with a sample monofilament so they could get a sense of what it feels like, and so that they understand that the stimulus is not painful. Participants wear a blindfold so that they can not see when the monofilament makes contact with their body, but do not wear headphones as the stimuli are inaudible. We then follow the protocol provided with the monofilament kit, mimicking how the procedure would be performed by clinicians [26]. The participants are instructed to say “yes” anytime they feel the filament touching them. If the participant does not feel the monofilament, we increase the monofilament evaluator size, otherwise we decrease the monofilament evaluator size. Once the participant does not feel a monofilament size (three times in a row for evaluator sizes 1.65 to 4.08 or one time in a row for evaluator sizes 4.17 to 6.65), we record the minimum evaluator size that they could feel as their monofilament threshold for that location. If the participant does not feel the starting monofilament, we increase the evaluator size until the subject can feel it and then record that evaluator size as their monofilament threshold for that location. As such, lower monofilament thresholds indicate better perception.
Sex assigned at birth | Male | 16 |
Female | 20 | |
Age | Older Adults (over 50) | 21 |
Younger Adults (18-50) | 15 | |
Older Mean (Range) | 66 (51,80) | |
Younger Mean (Range) | 27 (22,37) | |
@@
Black / of African Descent | 3 | |
Asian | 10 | |
White | 21 | |
@@
Handedness | Right | 32 |
Left | 2 | |
Ambidextrous | 2 | |
Footedness | Right | 31 |
Left | 2 | |
@@ II-B1 Participants
Twenty-one adult participants with no known history of diabetes or other disorders linked to peripheral neuropathy completed this study. Participant demographics are displayed in Table I. This study was approved by the Stanford University Institutional Review Board under Protocol 22514, and written consent was provided by all participants. Prior to completing the study, participants completed a pre-survey that inquired about demographic information as well hobbies or injuries that may impact touch sensitivity at the hands or feet.

### 25 Jan 2024 v2

@@ Abstract
Tactile perception plays an important role in activities of daily living, and it can be impaired in individuals with certain medical conditions. The most common tools used to assess tactile sensation, the Semmes-Weinstein monofilaments and the 128 Hz tuning fork, have poor repeatability and resolution. Long term, we aim to provide a repeatable, high-resolution testing platform that can be used to assess vibrotactile perception through smartphones without the need for an experimenter to be present to conduct the test. We present a smartphone-based vibration perception measurement platform and compare its performance to measurements from standard monofilament and tuning fork tests. We conducted a user study with 36 healthy adults in which we tested each tool on the hand, wrist, and foot, to assess how well our smartphone-based vibration perception thresholds (VPTs) detect known trends obtained from standard tests. The smartphone platform detected statistically significant changes in VPT between the index finger and foot and also between the feet of younger adults and older adults. Our smartphone-based VPT had a moderate correlation to tuning fork-based VPT. Our overarching objective is to develop an accessible smartphone-based platform that can eventually be used to measure disease progression and regression.
@@
Tactile perception, including vibrotactile perception, plays a critical role in enabling humans to perform various sensorimotor tasks such as object manipulation, navigation, and playing sports [1, 2, 3, 4]. We even use vibrotactile perception to balance during walking, a pervasive activity of daily living [5, 6, 7, 8, 9]. Many underlying health conditions and treatments including diabetes, chemotherapy, and direct injuries to the body can impair our tactile perception [10]. Given our reliance on tactile perception, deficits in perceiving tactile cues can have devastating consequences. Assessing tactile perception may help us understand disease progression and recovery, especially in response to treatment.
@@
Frequent use of vibration perception testing occurs during routine screenings of people experiencing peripheral neuropathy. The most common clinical diagnostics for peripheral neuropathy are the Semmes-Weinstein monofilament exam and the 128 Hz tuning fork exam [11]. Although these exams provide important information on tactile ability, they have limited measurement resolution. Tuning forks often provide inconsistent vibrations due to differences in how the clinician strikes the fork [12]. Monofilaments suffer from variations in force delivered due to variations in clinician application and overuse [13]. Also, when used in many clinical settings, both of these tools only measure a binary response (‘yes, can feel’ or ‘no, cannot feel’) to a single provided stimulus.
@@
Given the rise of smartphones with high-quality vibration actuators, there has been a rising interest in conducting mobile haptics experiments [14] [15]. We conducted preliminary work which showed that smartphone vibrations are more repeatable than tuning fork vibrations and that vibration perception threshold (VPT) could be reliably measured using a custom application employing a staircase algorithm [16]. Also in 2022, Torres et al. [17] characterized smartphone vibrations and found that smartphone-based VPTs correlate with monofilament-based pressure thresholds in the index finger. Several researchers have also worked toward validating the use of smartphone vibrations for diagnosing diabetic peripheral neuropathy. May et al. [18] showed that vibrations generated from a mobile phone could be used to detect diabetic peripheral neuropathy, and that the most accurate testing location was the first metatarsal head (a bony prominence in the big toe). Jasmin et al. [19] found that a vibration-based smartphone application has a moderate to strong correlation to tuning forks in classifying participants as experiencing neuropathy or not experiencing neuropathy, and that the interrater reliability using the smartphone application is high. While these studies are important steps towards developing an improved measurement tool, they suffer from various limitations including confounding factors in the study design, non-autonomous smartphone vibration perception data collection (reliance on an experimenter to physically conduct the exam), and the use of a binary measurement (yes/no) as opposed to a continuous, numerical value.
@@
In this work, we build upon our preliminary work by testing the feasibility of an application designed to measure smartphone-based vibration perception thresholds at different sites on the body (hand, wrist, and foot) in a healthy, age-diverse population. We also assess how the smartphone-based vibration perception thresholds correlate to tuning fork vibration perception thresholds and monofilament force perception thresholds at these locations. Our main goal is to confirm that smartphones can be used for vibration perception measurements and capture similar trends as currently used methods. In Section II, we describe our methods in characterizing, designing, and administering the smartphone-based VPT, performing the tuning fork and monofilament exams, and conducting a user study and accompanying analyses. We then discuss the study results in Section III and key takeaways and future work, such as testing our smartphone application on various patient populations and expanding this application to additional types of smartphones, in Section IV.
@@ II-A1 Smartphone VPT Exam
We developed an iOS application that controls Apple’s Core Haptics parameters (hapticIntensity and hapticSharpness) and autonomously implements a staircase algorithm (reversals = 8) to measure VPT. We chose to implement a one-up/one-down staircase method because it is a simple and fast way to calculate an absolute threshold. The one-up/one-down staircase method targets a performance level of 50% [20] [21], which is the standard performance level for absolute thresholds [20]. We chose to implement a one-up/one-down staircase instead of alternative one-up/two-down or one-up/three-down staircase methods that target higher performance levels because these alternatives require more trials and therefore more time to calculate a threshold; we chose to prioritize time as our eventual goal is to use this tool in a clinical setting where time is limited. Prior work by Yoshida and Kiernan, et al. [14] characterized the acceleration outputs at various locations on the phone; their results indicate that the output accelerations occur at a constant frequency of 230 Hz and that there is a positive, nonlinear relationship between the amplitude of the output acceleration and commanded hapticIntensity value when hapticSharpness is held constant at 1 and hapticIntensity is varied between 0.1 and 0.3. We further characterized this nonlinear relationship when hapticSharpness is held at a constant value of 1 and hapticIntensity is varied between 0 and 1 (Fig. 1).
@@
Despite this minor nonlinearity, we designed our staircase algorithm to output a continuous vibration for 0.1 seconds with the hapticSharpness set at 1.0 and the hapticIntensity varying with each step. This results in repeatable vibrations of varying amplitude with a constant frequency of 230 Hz, to which users can quickly and easily respond (Fig. 1). We chose to use hapticSharpness = 1 (230 Hz) since the Apple iPhone XS Max (our phone model) uses a linear resonant actuator that is tuned to operate at hapticSharpness = 1 (230 Hz). In addition to 230 Hz being the resonant frequency for this phone model’s linear resonant actuator, this frequency is also near the maximum sensitivity frequency of the Pacinian corpuscles, the mechanoreceptors largely responsible for our ability to feel high-frequency vibrations [22].
@@
The vibration amplitude starts small (hapticIntensity = 0.05) and increases with a step size of 0.05 until the vibration is detected by the user. At this point, the user says ‘yes’ to indicate detection, the spoken ‘yes’ is interpreted by the app using “Speech” (Apple’s voice recognition framework), and a reversal is recorded in the app. The vibration’s amplitude then decreases with a step size of 0.05 until the user can no longer detect the vibration (does not provide a spoken ‘yes’ response). Then, another reversal is recorded, the vibration’s amplitude increases again, and the staircasing of the hapticIntensity values continues until eight reversals are complete. Once complete, a CSV file containing the trial data is exported and stored in Google Firebase. The vibration perception threshold is calculated by averaging the hapticIntensity values at the eight reversal points (where the reversal point is the average of the value of the response that triggered a reversal and the value of the response prior to the reversal) as shown in Fig. 1. Time intervals between vibrations for each trial were randomly selected to reduce bias (ranging from 3-6 s), and responses had to occur within 2.5 seconds of the vibration in order to be recorded as a true positive response (as this is the upper end of haptic response times reported in literature [23, 24]). The smartphone VPT ranges from hapticIntensity of 0.05 (smallest vibration detected) to 1 (largest possible vibration output via Apple’s framework). As such, lower scores indicate better perception. If the largest vibrations output by the phone (vibrations generated by setting hapticIntensity to 1) were not felt three times in a row, the threshold for that trial is recorded as NaN because their threshold was outside the range of the smartphone-generated vibration stimuli.
@@
We measure VPT at six locations of interest: the index finger pad, the back of the index finger, the pinky finger pad, the dorsal wrist, the volar wrist, and the big toe pad (Fig. 2). These locations were chosen both for clinical relevance and ease of smartphone placement [25]. In Fig. 2B we also include the filtered vibration accelerations for the six configurations shown in Fig. 2A. For these measurements hapticIntensity was set to 0.25 and hapticSharpness was set to 1. We measure these waveforms using the setup described in detail in [16] and [14], with the modification that the phone was placed on a pillow while the body part was placed on the phone in the six different configurations. For each participant, smartphone-based VPT measurements are collected five times at each location, so that we can report the participant’s average smartphone-based VPT at each location.
@@
All smartphone-based measurements are collected on an Apple iPhone XS Max. Participants sit in a chair during the entire exam. For the finger and wrist locations, the phone is placed on a pillow that is placed on a desk. For the foot location, the phone is placed on a pillow, that is placed on the floor. Participants wore headphones playing their preferred music to prevent the use of auditory cues to identify vibrations. All but two participants listened to a Disney Hits playlist. Because the smartphone data collection component of the experiment took around an hour for most participants, we did not use white noise, which was found to make participants drowsy during piloting.
@@ II-A2 Tuning Fork Exam
We use a 128 Hz clinical tuning fork (CynaMed) to test the same six body parts as the smartphone (Fig. 2). Participants sit in a chair during the entire exam. For the finger and wrist locations, the hand/wrist is placed on a pillow that is placed on a desk. For the foot location, the foot is placed on a pillow, that is placed on a coffee table. The experimenter strikes the tines of the tuning fork on her knee and then places the base of the tuning fork on the body of the participant. Prior to collection, participants are touched with a vibrating tuning fork so they could get a sense of what it feels like and so that they understand that the sensation is not painful. Participants wear a blindfold so that they can not see when the tuning fork makes contact with their body, but do not wear headphones as the sound of the tuning fork vibrations are not easily discernible. Participants are instructed to say “start” when they start feeling a vibration and “stop” when they no longer feel any vibrations. A digital watch displaying seconds is used to measure the amount of time that the participant feels the vibration. The experimenter notes the second when the participant says “start” and also the second when the participant says “stop” and then records the difference in seconds, mimicking methods used in clinical settings. As such, higher tuning fork perception times indicate better perception. At each body part, tuning fork vibration perception time is measured five times. The times are then averaged, so that each participant’s average perception time at each of the body parts is reported.
@@ II-A3 Monofilament Exam
We use a 20-piece Semmes-Weinstein monofilament set (Touch Test Sensory Evaluators, North Coast Medical, Inc.) to assess light force perception. Monofilaments range from 0.008 grams-force to 300 grams-force. To assess light force perception at each of the six body parts (Fig. 2), we begin with the monofilament deemed normal for that location (0.07 grams-force for hands and dorsal feet, 0.4 grams-force for the plantar feet) [26]. Participants sit in a chair during the entire exam. For the finger and wrist locations, the hand/wrist is placed on a pillow that is placed on a desk. For the foot location, the foot is placed on a pillow, that is placed on a coffee table. Prior to collection, participants were touched with a sample monofilament so they could get a sense of what it feels like, and so that they understand that the stimulus is not painful. Participants wear a blindfold so that they can not see when the monofilament makes contact with their body, but do not wear headphones as the stimuli are inaudible. We then follow the protocol provided with the monofilament kit, mimicking how the procedure would be performed by clinicians [26]. The participants are instructed to say “yes” anytime they feel the monofilament touching them. If the participant does not feel the monofilament, we increase the monofilament evaluator size, otherwise we decrease the monofilament evaluator size. Once the participant does not feel a monofilament size (after three touches with the 0.008 grams-force to 1.000 grams-force monofilaments or one touch with the 1.400 grams-force to 300 grams-force monofilaments), we record the minimum evaluator size that they could feel as their monofilament threshold for that location. If the participant does not feel the starting monofilament, we increase the evaluator size until the subject can feel it and then record that evaluator size as their monofilament threshold for that location. As such, lower monofilament thresholds indicate better perception.
Sex assigned at birth | Male | 12 |
Female | 16 | |
Age | Older Adults (over 50) | 15 |
Younger Adults (18-50) | 13 | |
Older Mean (Range) | 67 (55,80) | |
Younger Mean (Range) | 28 (22,37) | |
@@
Black / of African Descent | 2 | |
Asian | 6 | |
White | 18 | |
@@
Handedness | Right | 26 |
Left | 1 | |
Ambidextrous | 1 | |
Footedness | Right | 25 |
Left | 1 | |
@@ II-B1 Participants
Thirty-six adult participants with no known history of diabetes or other disorders linked to peripheral neuropathy completed this study. Participant demographics are displayed in Table I. This study was approved by the Stanford University Institutional Review Board under Protocol 22514, and written consent was provided by all participants. Prior to completing the study, participants completed a pre-survey that inquired about demographic information as well as hobbies or injuries that may impact touch sensitivity at the hands or feet. Hand and foot dominance (defined as the foot one would use to kick a ball) were also inquired about via the survey.